OpenStudy (anonymous):

Find the standard matrix for the stated composition in R^2. a) A rotation of 90, followed by a reflection about the line y=x. b) An orthogonal projection on the y-axis, followed by a contraction with factor k=1/2. c) A reflection about the x-axis, followed by a dilation with factor k=3.

OpenStudy (loser66):

what is the matrix of rotation of 90?

OpenStudy (loser66):

that is \[\left[\begin{matrix}0&-1\\1&0\end{matrix}\right]\]

OpenStudy (anonymous):

it goes from \[\left[\begin{matrix}\cos \theta & -\sin \theta \\ \sin \theta & \cos \theta\end{matrix}\right]\] to\[\left[\begin{matrix}\sin \theta & -\cos \theta \\ \cos \theta & \sin \theta\end{matrix}\right]\]

OpenStudy (loser66):

now, the matrix to reflect about the line y =x is \[\left[\begin{matrix}0&1\\1&0\end{matrix}\right]\]

OpenStudy (loser66):

the same

OpenStudy (loser66):

but I use numbers to easily times them together

OpenStudy (loser66):

for a) rotation first, then reflection about y=x, so that the second matrix is written first, then the rotation matrix, that is \(\left[\begin{matrix}0&1\\1&0\end{matrix}\right]\)*\(\left[\begin{matrix}0&-1\\1&0\end{matrix}\right]\)

OpenStudy (loser66):

now, you compute the result

OpenStudy (anonymous):

oh, thanks, i had them reversed at first....

OpenStudy (loser66):

:)

OpenStudy (anonymous):

can you assist with part b?